OpenStudy (anonymous):

Let A, B, C be the angles of a triangle, so that A + B + C = pi. Show that cos(A+B) = -cosC

OpenStudy (gorv):

subtract C from both side ..u will get

OpenStudy (gorv):

\[\huge A+B=\pi-c\]

OpenStudy (gorv):

now take cosine both side

OpenStudy (gorv):

\[\huge \cos (\pi-C) =-cosC\]

OpenStudy (ikram002p):

cos(A+B)+cos(C)=0 cos(A+B)+cos(c)=0 2 cos ((A+B+C)/2) cos ( (A+B-C)/2)=0 cos ((A+B+C)/2) = cos(pi/2) =0 done :)